# Evaluation of Audio Compression Codecs

### v1

@@ 4.2 Perceptual Quality
Listening to the song, the guitar solo sounds slightly to the right, so the stereo image of this session should lean to the right of the center as well. All the sound-field visualizations in figures 4b to 4g have their clusters lean to the correct direction. However, the clusters in figures 4c to 4g, representing audio encoded with lossy compression techniques, significantly deviate from the uncompressed sound field in figure 4a. The sound-field clusters are too sparse for figures 4c to 4d using MP3 encoding, while the clustering for AAC CBR 256 kbps in figure 4e is too condensed in the center. However, both AAC VBR level 5 (cf. figure 4f) and Vorbis level 7 (cf. figure 4g) sound-field clusters are very similar to the uncompressed audio sound field in figure 4a.
Finally, advanced PEAQ ODG scores and totalNMR were calculated based on GstPEAQ [Holters:GSTPEAQ:2015] and [peaqbGitHub], respectively, to quantify the impact of compressing digital audio signals on perceptual quality. Table 5 provides advanced PEAQ ODG scores and totalNMR values for all tested samples in table 3. Advanced PEAQ ODG score range is 0 to -4 table 2 and the lower totalNMR the better.
Track Codec Setting Adv. ODG totalNMR
Dreams Fleetwood Mac FLAC level 6 0.2 -131.08
MP3 320 kbps -0.1 -0.77
MP3 128 kbps -0.1 -1.07
AAC 256 kbps -0.1 0.15
AAC level 5 -0.1 0.15
Vorbis level 7 -0.1 -16.75
La mer II. Jeux de Vagues Claude Debussy FLAC level 6 0.2 -121.33
MP3 320 kbps -2.0 -2.04
MP3 128 kbps -2.4 -2.4
AAC 256 kbps -1.4 -1.03
AAC level 5 -1.4 -1.03
Vorbis level 7 -0.1 -17.56
Foreigner Pallbearer FLAC level 6 0.2 -126.49
MP3 320 kbps -0.6 -3.04
MP3 128 kbps -0.2 -3.33
AAC 256 kbps -0.2 -2.35
AAC level 5 -0.2 -2.34
Vorbis level 7 -0.2 -17.28
September Earth, Wind & Fire FLAC level 6 0.2 -135.91
MP3 320 kbps -0.2 -1.14
MP3 128 kbps -3.5 -1.43
AAC 256 kbps -0.2 -0.46
AAC level 5 -0.2 -0.46
Vorbis level 7 -0.1 -16.74
Table 5: Advanced PEAQ ODG score and totalNMR for all tested samples in table 3.
Table 6 presents the average advanced PEAQ ODG scores and average totalNMR for all tested audio codecs at their most common quality settings.
Codec Setting Adv. ODG totalNMR
FLAC level 6 0.2 -130.11
MP3 CBR 320 kbps -0.2 -1.34
MP3 CBR 128 kbps -0.4 -1.72
AAC CBR 256 kbps -0.4 -0.08
AAC VBR level 5 -0.4 -0.08
Vorbis VBR level 7 -0.1 -16.74
Table 6: Average advanced PEAQ ODG scores and average totalNMR for all tested audio codecs. While all audio codecs score near 0 on advanced PEAQ ODG, FLAC’s average totalNMR is significantly lower than the average totalNMR of all lossy codecs, with Vorbis comes in second.
All audio codecs score near 0 on average advanced PEAQ ODG score. However, FLAC’s average totalNMR is the lowest. Vorbis’s average totalNMR is second lowest but there is a significant gap between FLAC and Vorbis. The remaining average totalNMR are -1 to 0.
@@ 4.3 Machine-Learning Based Codec
In recent years, a trend emerged to apply machine-learning techniques into remastering and compressing audio. One proposed audio compression model employs a residual vector quantized generative adversarial network (RVQGAN) [Kumar:HFACIRVQGAN:2023]. RVQGAN works by encoding the audio signal into a lower-dimensional presentation followed by converting this presentation into discrete values. We employed RVQGAN’s prototype implementation from its [DACRVQGANGithubRepo] and encoded the track Iron Man by Black Sabbath.
Table 7 presents RVQGAN’s average compression ratio, average encoding speed and average decoding speed together with traditional audio codecs for comparison.
Codec Setting Comp. Ratio Enc. Speed Dec. Speed
@@
Table 7: Average audio-codec compression as percent ratio and encoding as well as decoding speeds in samples per microsecond. Values for lossless and lossy encoding are the same as in table 4 plus the results for RVQGAN. Best performance in each category has been marked.
In its current state RVQGAN does improve on the compression ratio by reducing storage size to almost 1% of the original size. However, encoding and decoding speed are nearly one to two magnitudes slower than any of the traditional audio codecs.
Figures 5 to 7 illustrate the spectrum, spectrogram, and sound-field visualization of an RVQGAN encoded audio signal compared to the original uncompressed audio signal.
@@
Despite achieving an impressive compression ratio of around 98% of the uncompressed file size, perceptual quality of RVQGAN encoded audio is questionable. Currently, RVQGAN exhibits worse perceptual quality than all traditional audio codecs. RVQGAN’s average ODG score and average totalNMR in table 8 supports this evaluation.
Codec Setting Adv. ODG totalNMR
FLAC level 6 0.2 -130.11
MP3 CBR 320 kbps -0.2 -1.34
MP3 CBR 128 kbps -0.4 -1.72
AAC CBR 256 kbps -0.4 -0.08
AAC VBR level 5 -0.4 -0.08
Vorbis VBR level 7 -0.1 -16.74
RVQGAN -3.7 -0.69
Table 8: Average advanced PEAQ ODG scores and average totalNMR for all tested audio codecs. This is table 6 with the additional advanced PEAQ ODG score and totalNMR for RVQGAN.
@@ 5.1 Audio-Codec Performance
In our audio-codec performance tests (cf. table 4), FLAC exhibited the lowest compression ratio of all tested audio codecs, meaning that lossless audio files will take up more storage space than files encoded using a lossy compression technique. On the other hand, FLAC also achieved the highest encoding speed. These two average values for FLAC are expected because lossless compression techniques preserve most data in an audio file while only removing redundancies and repeated patterns.
All lossy audio codecs have similar compression ratio, with MP3 (128 kbps) achieving the highest compression ratio. Both MP3 quality settings exhibit lower encoding speeds than other lossy audio codecs. On the other hand, they show significantly higher decoding speeds than other lossy audio codecs with MP3 at 128 kbps being the fastest to decode. It can be argued that MP3 is a strong contender for the best lossy audio codec due to its high decoding speeds, making it the ideal choice for audio playback on low-power devices.
Among lossy audio encoders, Vorbis exhibits above average compression ratios, the fastest encoding speeds, and decent decoding speeds. In our audio-codec performance evaluation, Vorbis is the runner up for lossy audio-codecs, providing decent compression ratios and fastest lossy encoding speeds.
@@ 5.2 Perceptual Quality
Our perceptual quality evaluation shows that lossy compression algorithms sacrifice high-frequency aural details to reduce storage size. Audio encoding using psycho-acoustic principles determines these region are mostly inaudible to the average listener and the audio data for these regions can simply be removed. The differing graphs in figure 2 for all lossy audio codecs lies in how each codec handles loudness reduction of the high-frequency region.
@@
Loudness reduction or removal of high frequencies reduces the harmonic richness, the details and clarity, and the spatial depth [Truax:THFAE:2020] that allow digital music to exhibit high fidelity audio. [Vickers:TLWBSR:2010], mainly reporting on the loudness war in music production, also shows that the average human listener can perceive aural quality very differently depending on the loudness alteration in a music recording. Additionally, loudness reduction and high-frequency removals affect the stereo image of the encoded audio reducing their aural accuracy.
@@
The APEAQ ODG scores and totalNMR values we collected from all tested audio compression codecs (cf. table 6) support our findings. On average, all traditional audio codecs - FLAC, MP3, AAC, and Vorbis - score near 0 on APEAQ ODG. This means that APEAQ evaluates all of these traditional audio codecs impairment level as imperceptible (cf. table 2). However, the noise level of lossy codecs is higher than noise level of lossless codecs. In particular, noise levels for MP3 and AAC are right at the threshold to their mask levels, meaning their noise levels are potentially perceptible. Remarkably, Vorbis is the only lossy audio-compression codec where the noise level is masked reasonably well, making it the winner of this category.
@@
While most of our assessment was conducted via a system of codec performance measurements, visualizations, and PEAQ tests, we have to address the importance of listening tests. PEAQ were built under [ITU:BS.1387-2:2023] to predict how an average human perceive audio quality [ITU:BS.1387-2:2023]. Trained listeners, under [ITU:BS.1534-3:2015] on conducting MUSHRA tests, need to meet certain requirements to be qualified as assessors [ITU:BS.1534-3:2015]. Thus, they can provide important and practical feedback about how humans perceive aural quality, which validate the result of the objective evaluation models, hence validate our study.
@@ 5.3 AI-based Audio Codecs
Regarding AI-based compression technique, it is known that there were several neural-network models that were proposed. These models include an end-to-end approach [Rim:DNNETE:2021], lightweight transformer approach [Défossez:HFNAC:2022], and a novel audio codec called APCodec [Ai:APCodec:2024]. We chose the RVQGAN model by [Kumar:HFACIRVQGAN:2023] because the authors claimed their model outperforms competing traditional audio compression algorithms.
Our tests indicate RVQGAN indeed exhibits the highest compression ratio compared with traditional audio codecs (cf. table 7). Storage size of an RVQGAN encoded audio signal is only around 1% to 2% of the size of the uncompressed signal, which is an impressive result. However, RVQGAN’s encoding and decoding speeds are very low compared to traditional audio codecs, performing one to two magnitudes slower. Additionally, in its current version, RVQGAN’s output is a unique format descript-audio-codec (.dac) that is not playback-supported by any audio player application at the time of our evaluation tests.
While the RVQGAN codec achieved the highest average compression ratio in our audio-codec performance tests, Our testing shows that RVQGAN’s perceptual quality is worse than traditional audio codecs. RVQGAN’s average advanced ODG score is -3.7 which is close to very annoying (cf. table 2), and RVQGAN’s average totalNMR indicates that its noise level is quite perceptible (cv. table 8). The AI-based approach to developing audio codecs shows promise in advancing efficiency but will require further refinements to provide acceptable hi-fi listening experiences to end users.

### v2

@@ 4.2 Perceptual Quality
Figure 4 shows example sound-field visualizations for stereo images and sound stages capturing the 3:00 – 3:30 min interval of Iron Man by Black Sabbath when Tony Iommi performs the guitar solo. Listening to the uncompressed audio the guitar solo appears slightly to the right, which leads to the expectation that the stereo image of this session should lean to the right of the center as well. As expected the sound field for the FLAC encoded audio in figure 4b is nearly indistinguishable from that of the uncompressed audio. All the sound-field visualizations in figures 4b to 4g have their clusters lean in the expected direction. However, the clusters in figures 4c to 4g, representing audio encoded with lossy compression techniques, significantly deviate from the uncompressed sound field in figure 4a. The sound-field clusters are too sparse for figures 4c to 4d using MP3 encoding, while the clustering for AAC CBR 256 kbps in figure 4e is too condensed in the center. However, sound-field clusters for both AAC VBR level 5 (cf. figure 4f) and Vorbis level 7 (cf. figure 4g) are very similar to the uncompressed audio sound field in figure 4a.
Track Codec Setting BPEAQ ODG 2f Score APEAQ ODG BPEAQ totalNMR
Dreams Fleetwood Mac FLAC level 6 0.2 100 0.2 -131.08
MP3 320 kbps -3.8 32 -0.1 -0.77
MP3 128 kbps -3.8 32 -0.1 -1.07
AAC 256 kbps -3.9 29 -0.1 0.15
AAC level 5 -3.9 29 -0.1 0.15
Vorbis level 7 0.0 100 -0.1 -16.75
La mer II. Jeux de Vagues Claude Debussy FLAC level 6 0.2 100 0.2 -121.33
MP3 320 kbps -3.1 60 -2.0 -2.04
MP3 128 kbps -3.5 62 -2.4 -2.40
AAC 256 kbps -3.6 56 -1.4 -1.03
AAC level 5 -3.8 65 -1.4 -1.03
Vorbis level 7 0.0 100 -0.1 -17.56
Foreigner Pallbearer FLAC level 6 0.2 100 0.2 -126.49
MP3 320 kbps -3.2 71 -0.6 -3.04
MP3 128 kbps -3.6 71 -0.2 -3.33
AAC 256 kbps -3.6 63 -0.2 -2.35
AAC level 5 -3.5 63 -0.2 -2.34
Vorbis level 7 0.0 100 -0.2 -17.28
September Earth, Wind & Fire FLAC level 6 0.2 100 0.2 -135.91
MP3 320 kbps -3.5 29 -0.2 -1.14
MP3 128 kbps -3.6 30 -3.5 -1.43
AAC 256 kbps -3.7 27 -0.2 -0.46
AAC level 5 -3.8 27 -0.2 -0.46
Vorbis level 7 0.0 100 -0.1 -16.74
Table 5: Scores for basic PEAQ ODG, 2f model, and advanced PEAQ ODG as well as basic PEAQ totalNMR values for all tested audio tracks in table 3.
Our dataset was then tested using multiple PEAQ experiments and overall results for a variety test tracks are shown in table 5. As expected, lossless FLAC achieves perfect scores under all metrics. This means FLAC’s perceptual quality is near-identical to the original audio’s perceptual quality. Both MP3 320 kbps and MP3 128 kbps, along with AAC 256 kbps and AAC VBR level 5, score near -4 on BPEAQ, indicating very perceptible impairment (cf. table 2). The 2f score of both MP3 codecs as well as both AAC codecs range from 27 (poor) to 71 (fair) indicating that, if these audio codecs were tested in MUSHRA, trained listeners would find these four audio codecs somewhat decent on some tracks and below average on other tracks.
The observed differences in 2f scores may partly reflect variations in the spectral and temporal characteristics of the source material. A pop song like Fleetwood Mac’s Dreams as well as material from the disco genre, e.g., Earth, Wind & Fire’s September, contain dense rhythmic instrumentation including prominent percussion and sustained mid–to-high frequency energy, which increase temporal modulation complexity and may amplify codec-induced modulation distortion. In contrast, a classical performance such as Debussy’s La mer exhibits more structured orchestral textures with comparatively smoother spectral evolution. Rock performances like Pallbearer’s Foreigner contain heavily distorted guitar signals whose broadband harmonic content can raise masking thresholds and partially conceal additional codec artifacts. These genre dependent signal properties may contribute to the better 2f scores observed for the latter tracks.
Using APEAQ, most encoded tracks show good ODG score of 0 and -1. The only exceptions are Debussy’s La mer, scoring at -2 for MP3 320 kbps and MP3 128 kbps, and Earth, Wind & Fire’s September, scoring worst at -3.5 for MP3 128 kbps, indicating perceptible impairment (cf. table 2). This indicates that the compression impairment in MP3 and AAC audio codecs may not be as severe as previously shown by BPEAQ on some tracks. However, audio impairment of Earth, Wind & Fire’s September encoded with MP3 128 kbps is rated in both BPEAQ and APEAQ as quite perceptible. totalNMR values for both MP3 audio-codec settings as well as both AAC audio-codec settings range from -3 to 0 meaning the noise to mask ratio is rather high and most noise from these encoded test tracks will be audible to listeners.
The most interesting case in table 5 is Vorbis VBR level 7. It is the only lossy audio codec that scores 0 on both BPEAQ ODG and APEAQ ODG as well as 100 on 2f score. This strongly indicates Vorbis’ audio impairment is imperceptible to listeners and the audio codec’s perceptual quality is very close to lossless FLAC perceptual quality. Vorbis’ totalNMR value is around -17, which means its noise level is less audible than the remaining tested audio codecs’ noise level.
Codec Setting BPEAQ ODG 2f Score APEAQ ODG BPEAQ totalNMR
FLAC level 6 0.2 100 0.2 -130.11
MP3 CBR 320 kbps -3.6 36 -0.2 -1.34
MP3 CBR 128 kbps -3.7 36 -0.4 -1.72
AAC CBR 256 kbps -3.7 32 -0.4 -0.08
AAC VBR level 5 -3.8 37 -0.4 -0.08
Vorbis VBR level 7 0.0 99 -0.1 -16.74
Table 6: Average scores for basic PEAQ ODG, 2f model, and advanced PEAQ ODG as well as PEAQ totalNMR values for all tested audio codecs (cf. table 6). Best performance in each category has been marked.
Table 6 presents the average BPEAQ ODG scores, average 2f scores, average APEAQ ODG scores as well as average BPEAQ totalNMRs for all tested audio codecs using the quality settings defined in table 6. BPEAQ ODG for FLAC is 0 and for Vorbis 0.2 on average, translating to perceptual quality (nearly) the same as the original audio signal. The remaining audio codecs’ BPEAQ ODG average close to -4, which means theses audio codecs’ impairment is classified as very annoying (cf. table 2). These partial results are supported as well by the average 2f-model scores, which indicate that FLAC and Vorbis would score 100 on the MUSHRA listening test while the remaining codecs perform at much lower values in the 30s. The average totalNMR also shows that noise in all lossy audio codecs is very noticeable except for Vorbis. The interesting case here is the APEAQ ODG score, which shows much better values for all audio codecs at near 0. This indicates that APEAQ regards all tested audio codecs impairment as imperceptible, which contradicts with all the BPEAQ ODG scores, 2f-model scores, and totalNMR values.
@@ 4.3 Machine-Learning Based Codec
In recent years, a trend emerged to apply machine-learning techniques into remastering and compressing audio. One proposed audio compression model employs a residual vector quantized generative adversarial network (RVQGAN) [undefw]. RVQGAN works by encoding the audio signal into a lower-dimensional presentation followed by conversion into discrete values. We employed RVQGAN’s prototype implementation from its [undefg] and encoded the Black Sabbath’s Iron Man with it.
Codec Setting Compression Ratio Encoding Speed Decoding Speed
@@
Table 7: Average audio-codec compression as percent ratio and encoding as well as decoding speeds in samples per microsecond. Values for lossless and lossy encoding are the same as in table 4 with an additional result row for RVQGAN. Best performance in each category has been marked.
Table 7 presents RVQGAN’s average compression ratio, average encoding speed, and average decoding speed in comparison with traditional audio codecs previously evaluated. In its current state, RVQGAN does improve on the compression ratio by reducing storage size to almost 1% of the original size. However, encoding and decoding speed are nearly one to two magnitudes slower than any of the traditional audio codecs.
@@
Figures 5 to 7 illustrate the spectrum, spectrogram, and sound-field visualization of an RVQGAN encoded audio signal compared to the original uncompressed audio signal. The spectrogram in figure 6 shows a cutoff in loudness at around 21 kHz. The spectrum in figure 5 shows the RVQGAN tested track exhibits first a slight loudness increase in the lower frequency range, which gradually decreases form around 10 kHz to -100 dB at 20 kHz and beyond. figure 7 shows the encoded audio-signal output pattern more smeared and less accurate compared to the uncompressed track’s original pattern. Overall, these visualizations indicate that RVQGAN encoded audio signals may share similar perceptual quality problems with lossy encoded audio.
Codec Setting BPEAQ ODG 2f Score APEAQ ODG BPEAQ totalNMR
FLAC level 6 0.2 100 0.2 -130.11
MP3 CBR 320 kbps -3.6 36 -0.2 -1.34
MP3 CBR 128 kbps -3.7 36 -0.4 -1.72
AAC CBR 256 kbps -3.7 32 -0.4 -0.08
AAC VBR level 5 -3.8 37 -0.4 -0.08
Vorbis VBR level 7 0.0 99 -0.1 -16.74
RVQGAN -2.7 37 -3.7 -0.69
Table 8: Average scores for basic PEAQ ODG, 2f model, and advanced PEAQ ODG as well as PEAQ totalNMR values for all tested audio codecs (cf. table 6) plus average scores for RVQGAN. Best performance in each category has been marked.
Despite achieving an impressive compression ratio of around 98% of the uncompressed file size, perceptual quality of RVQGAN encoded audio is questionable. Currently, RVQGAN exhibits worse perceptual quality than all traditional audio codecs. RVQGAN’s average evaluation scores in table 8 supports this evaluation.
@@ 5.1 Audio-Codec Performance
In our audio-codec performance tests (cf. table 4), FLAC exhibited the lowest compression ratio of all tested audio codecs meaning that lossless audio files will take up more storage space than files encoded using a lossy compression technique. On the other hand, FLAC achieved the highest encoding speed. These two average values for FLAC are expected because lossless compression techniques preserve most of the information in an audio signal while only compressing redundancies and repeated patterns.
All lossy audio codecs have similar compression ratios with MP3 128 kbps achieving the best results. Both MP3 quality settings, i.e. 128 kbps and 320 kbps, exhibit lower encoding speeds than other lossy audio codecs. On the other hand, these codecs show significantly higher decoding speeds than other lossy audio codecs with MP3 128 kbps exhibiting the fastest decoding speed. It can be argued that MP3 is a strong contender for the best lossy audio codec due to its high decoding speeds, making it the ideal choice for audio playback on low-power devices. This is because MP3 has lower computational complexity compared to modern codecs like AAC or Vorbis, which is due to its older more simple perceptual encoding model [undefc]. However, MP3’s encoding speed is constrained by its two-nested-iteration-loop design [undefc], which makes MP3’s encoding speed fall behind when comparing to AAC and Vorbis newer and more sophisticated encoding techniques.
Among lossy audio encoders, Vorbis exhibits above average compression ratios, the fastest encoding speeds, and decent decoding speeds. In our audio-codec performance evaluation, Vorbis is the runner up for lossy audio codecs, providing good compression ratios and fastest lossy encoding speeds.
@@ 5.2 Perceptual Quality
Our perceptual quality evaluation shows that, while lossy compression algorithms still try to preserve perceptual audio quality, they sacrifice high-frequency aural details to reduce storage size, thus reducing the quality of the audio signal. Because audio encoding using psycho-acoustic principles determines regions that are mostly inaudible to the average listener, audio data for these regions are simply removed. The spectra for all lossy audio codecs (cf. figure 2) clearly show how each lossy audio codec handles loudness reduction of high-frequency regions.
@@
Loudness reduction and removal of high frequencies reduces the harmonic richness, the details and clarity, and the spatial depth [undefah] that allow digital music to exhibit high-fidelity audio qualities. [undefai], mainly reporting on the loudness war in music production, also shows that the average human listener is able to perceive aural quality very different depending on the loudness alteration in a music recording. Additionally, loudness reduction and high-frequency removals negatively affect the stereo image of the encoded audio signal reducing aural accuracy (cf. figure 1c).
We employ multiple complementary PEAQ predictors, including basic PEAQ, advanced PEAQ, basic PEAQ’s totalNMR values, and the 2f model. Using a combination of not just ODG scores but also MOV scores allows to capture distinct psychoacoustic dimensions. Convergence across these metrics (cf. table 6) strengthens confidence that observed audio-codec artifacts align with listener perceptible phenomena, providing a reproducible computational approximation rather than a substitute for subjective evaluation.
@@
The ODG scores and the MOV scores collected from all tested audio compression codecs (cf. table 6) support these findings. On average, all traditional audio codecs, i.e. FLAC, MP3, AAC, and Vorbis, score near 0 on advanced PEAQ ODG. This means that advanced PEAQ classifies all of these traditional audio codecs’ impairment level as imperceptible (cf. table 2). However, the noise level of lossy audio codecs is higher than the noise level of lossless audio codecs. In particular, noise levels for MP3 and AAC are right at the threshold of their masking levels, meaning their noise levels are potentially perceptible. Remarkably, Vorbis is the only lossy audio-compression codec where the noise level is masked reasonably well, making it the winner in this category. The 2f-model scores support this as well with Vorbis’ achieving an average score of 99 while other lossy encoders’ average 2f-model scores are in the 30s.
The notable discrepancy between basic PEAQ ODG scores and advanced PEAQ ODG scores observed for encoded audio with MP3 and AAC (cf. table 6) warrants further discussion. While basic PEAQ yields an average ODG score of approximately -3.7 for both codecs, suggesting very annoying impairment, advanced PEAQ in contrast produces scores near 0, suggesting imperceptible impairment. This gap can be attributed to the difference in perceptual modeling depth between the two models. Basic PEAQ’s MOVs, which rely primarily on noise-to-mask ratio and loudness-based metrics [undefaf], can be rather rigid when evaluating audio artifacts above basic PEAQ’s masking threshold, even if not be perceptible for a human listener. Advanced PEAQ, by contrast, incorporates MOVs capturing modulation masking, harmonic coherence, and asymmetric distortion weighting [undefaf], which together better account for the structured, partially-masked nature of audio encoding artifacts.
The above findings suggests that basic PEAQ’s ODG score should be interpreted with caution when evaluating modern perceptual codecs. The magnitude of the gap between basic ODG scores to advanced ODG scores in itself may serve as an indicator of how well a codec’s artifact profile aligns with a given perceptual model’s assumptions. Our finding of a gap between basic PEAQ ODG scores and advanced PEAQ ODG scores also agrees with work by [undefj], which motivated conducting additional evaluation with the 2f model and basic PEAQ’s totalNMR MOV. Together these four metrics allow a comprehensive and objective comparison of an audio codec’s perceptual quality characteristics.
@@
While most of our assessment was conducted via a system of audio codec performance measurements, visualizations, and various PEAQ-related tests, we have to address the importance of subjective listening tests. PEAQ variants were built under [undefr] to predict how an average human listener perceives audio quality [undefr]. Trained listeners, under [undefs] conducting MUSHRA tests, are obligated to meet certain (minimum) requirements to qualify as assessors [undefs]. Thus, trained listeners may be able to provide additional feedback about how humans perceive aural quality, which would be useful to validate the results of the objective evaluation models and, hence, validate and improve our study and approach.
@@ 5.3 AI-based Audio Codecs
Our tests indicate RVQGAN [undefw] indeed exhibits the highest compression ratio compared with traditional audio codecs (cf. table 7). Storage size of an RVQGAN encoded audio signal is only around 1% to 2% of the size of the uncompressed signal, which is an impressive result. However, RVQGAN’s encoding and decoding speeds are very low compared to traditional audio codecs, performing one to two magnitudes slower. Additionally, in its current version RVQGAN’s output is a unique file format, descript-audio-codec (.dac) [undefg], which is not playback-supported by any audio player application at the time of our evaluation tests.
While the RVQGAN codec achieved the highest average compression ratio in our audio-codec performance tests, our testing shows that RVQGAN’s perceptual quality is at the worse end of traditional audio codecs’ perceptual quality. RVQGAN’s average advanced ODG score is -3.7, which is close to very annoying (cf. table 2), and RVQGAN’s average totalNMR value indicates that its noise level is quite perceptible (cf. table 8). AI-based approaches to developing or discovering audio codecs shows promise in advancing efficiency but will require further refinements to provide acceptable hi-fi listening experiences to end users.